# Bridging Simplicity and Sophistication using GLinear: A Novel Architecture for Enhanced Time Series Prediction

### v2

@@ Abstract
Time Series Forecasting (TSF) is an important application across many fields. There is a debate about whether Transformers, despite being good at understanding long sequences, struggle with preserving temporal relationships in time series data. Recent research suggests that simpler linear models might outperform or at least provide competitive performance compared to complex Transformer-based models for TSF tasks. In this paper, we propose a novel data-efficient architecture, GLinear, for multivariate TSF that exploits periodic patterns to provide better accuracy. It also provides better prediction accuracy by using a smaller amount of historical data compared to other state-of-the-art linear predictors. Four different datasets (ETTh1, Electricity, Traffic, and Weather) are used to evaluate the performance of the proposed predictor. A performance comparison with state-of-the-art linear architectures (such as NLinear, DLinear, and RLinear) and transformer-based time series predictor (Autoformer) shows that the GLinear, despite being parametrically efficient, significantly outperforms the existing architectures in most cases of multivariate TSF. We hope that the proposed GLinear opens new fronts of research and development of simpler and more sophisticated architectures for data and computationally efficient time-series analysis. The source code is publicly available on GitHub.
@@ I Introduction
Accurate forecasting has become increasingly valuable in today’s data-driven world, where computational intelligence is key to automated decision-making [10, 18]. Time Series Forecasting (TSF) tasks have various applications that impact diverse fields such as finance, healthcare, supply chain management, and climate science [12]. The ability to predict future trends based on historical data not only enhances decision-making processes, but also drives innovation and operational efficiency.
@@
The emergence of advanced and hybrid methods that use Deep Learning (DL), have revolutionized predictive modeling by enabling the extraction of complex patterns within the data [34]. Recurrent Neural Networks (RNNs), Long Short Term Memory (LSTM), and Transformer-based architectures have shown promising results in medium- and long-term forecasting without rigid assumptions [21, 1, 22]. Despite their sophistication, recent studies indicate that linear models can capture periodic patterns and provide competitive performance in certain contexts along with computational efficiency [1, 2].
Not all time-series data are suitable for precise predictions, particularly when it comes to long-term forecasting, which becomes especially difficult in chaotic systems [2]. Long-term forecasting has been shown to be most feasible for time series data when data exhibits clear trends and periodic patterns  [2, 1]. This brings up a question: How can we effectively integrate the simplicity of linear models with sophisticated techniques for capturing complex underlying patterns to further enhance medium- and long-term TSF?
@@
Inspired from long TSF linear models (NLinear, DLinear and RLinear); we propose a novel data-efficient architecture, GLinear. GLinear is a simple model that does not have any complex components, functions, or blocks (like self-attention schemes, positional encoding blocks, etc.) like previously mentioned linear models. It has capabilities of enhanced forecasting performance while maintaining simplicity. Furthermore, GLinear focuses on data efficiency and demonstrates the potential to perform robust forecasting without relying on extensive historical datasets, which is a common limitation of other TSF models. Our contributions in this paper are outlined as follows:
  • We propose a novel architecture, GLinear that can deliver highly accurate TSF by leveraging periodic patterns, making it a promising solution for diverse forecasting applications.
@@
  • We explore GLinear’s applicability across diverse sectors, assessing its impact on forecasting performance with varying data input length and the prediction horizons.
The remainder of the paper is organized as follows: Section II presents related work on TSF using state-of-the-art transformer-based models and linear predictors along with their limitations. Section III presents the architecture of different linear models to provide a better understanding of the proposed method. Section IV presents the proposed GLinear model. Section V presents the detail of experimental setup including used datasets, implementation details and evaluation metrics. Different experiments and their results are presented in Section VI. Finally, Section VII concludes our research with takeaways and future directions.
@@ II-A State-of-the-Art Transformers
Transformer architectures have demonstrated remarkable potential in time series forecasting by effectively capturing long-range dependencies, a critical aspect often overlooked by traditional methods. However, adapting Transformers for time series requires addressing inherent challenges such as computational complexity and the lack of inherent inductive biases for sequential data [23, 27]. This efficiency bottleneck is addressed with the Informer [24] model with the introduction of ProbSparse attention, which reduces complexity from $O(L^{2})$ to $O(L*logL)$ and enables efficient processing of long sequences [24]. They also employed a generative decoder, predicting long sequences in a single forward pass. Another version of the transformer model, Autoformer [25], was proposed to tackle the same complexity issue by replacing an auto-correlation mechanism with the dot product attention to efficiently extract dominant periods in the time series. Their approach proved particularly effective for long-term forecasting on datasets like ETT and Electricity Transformer [25]. Furthermore, Wu et al. [26] incorporated time-specific inductive biases in their approach. Their proposed model, TimesNet, was introduced to treat time series as images and leverage 2D convolution operations across multiple time scales to capture intra- and inter-variable relationships, achieving state-of-the-art results on various long-term forecasting benchmarks [26].
The flexibility of Transformer-based models has facilitated their application across diverse forecasting horizons and domains. For short- and medium-term forecasting, adaptations focusing on computational efficiency and local pattern extraction have been explored. For instance, FEDformer [27] proposed frequency-enhanced attention and a mixture of expert decoders to capture both global and local patterns efficiently. This approach has shown promising results in short-term load forecasting and other applications where capturing high-frequency components is crucial. For long-term forecasting, the ability of Transformers to model long-range dependencies becomes paramount. TimesNet has demonstrated remarkable performance in this domain [26]. Furthermore, some recent researches [28, 29] have utilized external factors and contextual information into Transformer models. Such as integrating weather data or economic indicators to improve forecasting accuracy in domains like energy consumption and financial markets. Additionally, probabilistic forecasting using Transformers is gaining traction, providing not only point predictions but also uncertainty quantification, which is essential for risk management in various applications [28, 29].
@@ II-B State-of-the-Art Linear Predictors
One prominent research direction focuses on enhancing classical linear methods to better capture complex temporal dynamics. Traditional methods like Autoregressive (AR) [35] models and their variants, while computationally efficient, often struggle with non-linear patterns and long-range dependencies. Zheng et al. [33] introduced a dynamic regression technique that allows the model coefficients to vary over time using adaptive filtering. This approach dynamically adjusts model parameters based on incoming data, improving the adaptability of these models to changing time series characteristics [33]. Other techniques utilizing Kalman filtering within a linear framework have demonstrated effectiveness in tracking evolving trends and seasonality [36]. Furthermore, it has been shown that the use of sparse linear models, such as LASSO regression, which select only the most relevant past observations for prediction, enhances both efficiency and interpretability [31]. These advancements aim to maximize the performance of established linear frameworks by integrating sophisticated techniques to mitigate their inherent limitations. Another research direction involves developing novel linear architectures specifically tailored for time series data. This includes models like DLinear which decomposes the time series into trend and seasonal components and models them with simple linear layers, achieving surprisingly strong performance on long-term forecasting tasks [1]. Similarly, N-Linear proposes a simple neural network with a single linear layer for forecasting, demonstrating competitive results while drastically reducing computational complexity [1].
Although linear predictors like NLinear, DLinear, and RLinear have achieved competitive accuracy with drastically reduced computational overhead; these models still exhibit some limitations. Such as struggling with capturing complex non-linear patterns or failing to effectively model specific time series characteristics, such as strong seasonality or abrupt changes in trend. Furthermore, these models also rely heavily on extensive historical data to achieve high prediction accuracy. This dependency can limit their effectiveness in scenarios with limited data availability, such as newly established systems or rapidly changing environments. These issues motivate the development of novel architectures and improvements in linear models, which aim to address these shortcomings.
GLinear achieves superior results by integrating two components: (1) a non-linear Gaussian Error Linear Unit (GeLU)-based transformation layer to capture intricate patterns, and (2) Reversible Instance Normalization (RevIN) to standardize data distributions across instances, ensuring consistent performance and adaptability across diverse datasets. This approach provides a more comprehensive and efficient solution for TSF.
@@ III Architecture of Different Linear Models
The input of a time series predictor is the $L$ timesteps of past input samples (also referred to as the lookup window or input sequence length). A predictor uses this input to predict $T$ timesteps of future values (also referred to as output horizon or prediction length). The details of different linear predictors are given below:
Figure 1: Architecture of LTSF-Linear.
Figure 2: (a) The architecture of Nlinear predictor, (b) The architecture of DLinear predictor. The illustration of the architectures is created based on the description in the original paper [1].
@@ III-A Linear Predictor (LTSF-Linear)
The first model is a LTSF-Linear predictor [1] where LTSF stands for long-term TSF. The architecture of the linear predictor can be visualized from Figure 1. This predictor is composed of a single fully connected linear layer, or dense layer. LSTF-Linear predictor does not model any spatial correlations. A single temporal linear layer directly regresses historical time series for future prediction via a weighted sum operation as follows:
$T=WL+b$ (1)
where $L$ refers to past input samples depending on defined input sequence length (can also be denoted as $X$), $T$ refers to predicted future samples depending on required prediction length (can also be denoted as $\hat{X}$), $W$ is weight matrix $W\in R^{T\times L}$ and $b$ is additive bias.
@@ III-C DLinear Predictor (Decomposition-based Linear Model)
DLinear [1] is a combination of a decomposition scheme used in Autoformer and FEDformer with linear layers. It first decomposes raw data input into a trend component by a moving average kernel and a remainder (seasonal) component as shown in Figure 2 (b). Then, two one-layer linear layers are applied to each component, and the two features are sum up to get the final prediction. By explicitly handling trend, DLinear enhances the performance of a vanilla linear when there is a clear trend in the data.
@@ IV Methodology
Existing linear models use different simple operations like normalization and decomposition. It is worth noting that with the involvement of simpler mathematical operations, it is possible to build powerful linear predictors with some variations of activation functions to extract meaningful results for the required task [3]. Keeping these in mind, a new Gaussian-based linear predictor, GLinear, is proposed.
@@
Figure 4 shows the architecture of the GLinear predictor that is composed of two fully-connected layers of the same input size having a GeLU nonlinearity in-between them. Different configurations of input and layers sizes are tested to lead to this final architecture. GeLU [5] is a non-linear activation function defined as follow:
$\operatorname{GELU}(x)=x\cdot\Phi(x)$ (2)
where $\Phi(x)$ refers to the cumulative distribution function (defined in (3)) of the standard normal distribution with an error function (erf).
$\Phi(x)=x\cdot\frac{1}{2}\left(1+\text{erf}\left(\frac{x}{\sqrt{2}}\right)\right)$ (3)
@@
Furthermore, RevIN is applied to the input and output layers of the GLinear model [4]. This normalization layer transforms the original data distribution into a mean-centred distribution, where the distribution discrepancy between different instances is reduced. This normalized data is then applied as new input to the used predictor, and then the final output is denormalized at the last step to provide the final prediction. RevIN can be used with any existing model. It does not add any significant overhead in training time due to low computational complexity of this normalization scheme.
@@
The time-series data usually suffer from distribution shift problem due to changes in its mean and variance over time that can degrade the performance of time series predictors. RevIN utilizes learnable affine transformation to remove and restore the statistical information of a time-series instance that can be helpful to handle datasets with distribution shift problem.
Some features of GLinear model are:

### v4

@@ Abstract
Time Series Forecasting (TSF) is an important application across many fields. There is a debate about whether Transformers, despite being good at understanding long sequences, struggle with preserving temporal relationships in time series data. Recent research suggests that simpler linear models might outperform or at least provide competitive performance compared to complex Transformer-based models for TSF tasks. In this paper, we propose a novel data-efficient architecture, Gaussian-activated Linear model (GLinear), for multivariate TSF that exploits periodic patterns to provide better accuracy. It achieves higher prediction accuracy while requiring less historical data than other state-of-the-art linear predictors. Four different datasets (ETTh1, Electricity, Traffic, and Weather) are used to evaluate the performance of the proposed predictor. A performance comparison with state-of-the-art linear architectures (such as NLinear, DLinear, and RLinear) and transformer-based time series predictors (Autoformer) shows that the GLinear, despite being data efficient, outperforms the existing architectures in most cases of multivariate TSF while being competitive in others. We hope that the proposed GLinear model opens new fronts of research and development of simpler and more sophisticated architectures for data and computationally efficient time-series analysis. The source code is publicly available on GitHub.
@@ I Introduction
Accurate forecasting has become increasingly valuable in today’s data-driven world, where computational intelligence plays a key role in automated decision-making [1, 2]. Time Series Forecasting (TSF) tasks have various applications that impact diverse fields such as finance, healthcare, supply chain management, and climate science [3]. The ability to predict future trends based on historical data not only enhances decision-making processes but also drives innovation and operational efficiency.
@@
The emergence of advanced and hybrid methods that use Deep Learning (DL) has revolutionized predictive modeling by enabling the extraction of complex patterns within the data [10]. Recurrent Neural Networks (RNNs), Long Short-Term Memory (LSTM), and Transformer-based architectures have shown promising results in medium- and long-term forecasting without rigid assumptions [11, 12, 13]. Despite their sophistication, recent studies indicate that linear models can capture periodic patterns and provide competitive performance in certain contexts, along with computational efficiency [12, 14].
Not all time-series data are suitable for precise predictions, particularly when it comes to long-term forecasting, which becomes especially difficult in chaotic systems [14]. Long-term forecasting is most feasible for time series data when data exhibits clear trends and periodic patterns  [14, 12]. This brings up a question: How can we effectively integrate the simplicity of linear models with sophisticated techniques for capturing complex underlying patterns to further enhance medium- and long-term TSF?
@@
Inspired by long TSF linear models (NLinear [12], DLinear [12], and RLinear [14]), we propose a novel data-efficient architecture, Gaussian-activated Linear (GLinear) model. GLinear is a simple model that does not have any complex components, functions, or blocks (like self-attention schemes, positional encoding blocks, etc.), like previously mentioned linear models. It has capabilities of enhanced forecasting performance while maintaining simplicity. Furthermore, GLinear focuses on data efficiency and demonstrates the potential to perform robust forecasting without relying on extensive historical datasets, which is a common limitation of other TSF models. Our contributions in this paper are outlined as follows:
  • We propose a novel architecture, GLinear, that can deliver highly accurate TSF by leveraging periodic patterns, making it a promising solution for diverse forecasting applications.
@@
  • We explore GLinear’s applicability across diverse sectors, assessing its impact on forecasting performance with varying data input length and prediction horizons.
The remainder of the paper is organized as follows: Section II presents related work on TSF using state-of-the-art transformer-based models and linear predictors, along with their limitations. Section III presents the architecture of different linear models to provide a better understanding of the proposed method. Section IV presents the proposed GLinear model. Section V presents the details of the experimental setup, including the used datasets, implementation details, and evaluation metrics. Different experiments and their results are presented in Section VI. Finally, Section VII concludes our research with takeaways and future directions.
@@ II-A State-of-the-Art Transformers
Transformer architectures have demonstrated remarkable potential in different DL tasks by effectively capturing long-range dependencies, a critical aspect often overlooked by traditional methods [19, 20, 21]. However, adapting Transformers for time series requires addressing inherent challenges such as computational complexity and the lack of inherent inductive biases for sequential data [15, 22]. This efficiency bottleneck is addressed with the Informer [16] model with the introduction of ProbSparse attention, which reduces complexity from $O(L^{2})$ to O(LlogL)O(L*logL) and enables efficient processing of long sequences [16]. They also employed a generative decoder, predicting long sequences in a single forward pass. Another version of the transformer model, Autoformer [17], was proposed to tackle the same complexity issue by replacing an auto-correlation mechanism with the dot product attention to efficiently extract dominant periods in the time series. Their approach proved particularly effective for long-term forecasting on datasets like ETTh and Electricity Transformer [17]. Furthermore, Wu et al. [23] incorporated time-specific inductive biases in their approach. Their proposed model, TimesNet, was introduced to treat time series as images and leverage 2D convolution operations across multiple time scales to capture intra- and inter-variable relationships, achieving state-of-the-art results on various long-term forecasting benchmarks  [23].
Figure 1: The architecture of LTSF-Linear.
The flexibility of Transformer-based models has facilitated their application across diverse forecasting horizons and domains. For short- and medium-term forecasting, adaptations focusing on computational efficiency and local pattern extraction have been explored. For instance, FEDformer [22] proposed frequency-enhanced attention and a mixture of expert decoders to capture both global and local patterns efficiently. This approach has shown promising results in short-term load forecasting and other applications where capturing high-frequency components is crucial. For long-term forecasting, the ability of Transformers to model long-range dependencies becomes paramount. Times-Net has demonstrated remarkable performance in this domain [23]. Furthermore, some recent research [24, 25] has incorporated external factors and contextual information into Transformer models. Such as integrating weather data or economic indicators to improve forecasting accuracy in domains like energy consumption and financial markets. Additionally, probabilistic forecasting using Transformers is gaining traction, providing not only point predictions but also uncertainty quantification, which is essential for risk management in various applications [24, 25].
@@ II-B State-of-the-Art Linear Predictors
Figure 2: (a) The architecture of the NLinear predictor, (b) The architecture of the DLinear predictor. The illustration of the architecture is created based on the description in the original paper [12].
One prominent research direction focuses on enhancing classical linear methods to better capture complex temporal dynamics. Traditional methods like auto-regressive (AR) [28] models and their variants, while computationally efficient, often struggle with non-linear patterns and long-range dependencies. Zheng et al. [29] introduced a dynamic regression technique that allows the model coefficients to vary over time using adaptive filtering. This approach dynamically adjusts model parameters based on incoming data, improving the adaptability of these models to changing time series characteristics [29]. Other techniques utilizing Kalman filtering within a linear framework have demonstrated effectiveness in tracking evolving trends and seasonality [30]. Furthermore, it has been shown that the use of sparse linear models, such as LASSO regression, which select only the most relevant past observations for prediction, enhances both efficiency and interpretability [31]. These advancements aim to maximize the performance of established linear frameworks by integrating sophisticated techniques to mitigate their inherent limitations. Another research direction involves developing novel linear architectures specifically tailored for time series data. This includes models like DLinear [12], which decomposes the time series into trend and seasonal components and models them with simple linear layers, achieving surprisingly strong performance on long-term forecasting tasks [12]. Similarly, NLinear [12] proposes a simple neural network with a single linear layer for forecasting, demonstrating competitive results while drastically reducing computational complexity.
Although linear predictors like NLinear [12], DLinear [12], and RLinear [14] have achieved competitive accuracy with drastically reduced computational overhead, these models still exhibit some limitations. Such as struggling with capturing complex non-linear patterns or failing to effectively model specific time series characteristics, such as strong seasonality or abrupt changes in trend. Furthermore, these models also rely heavily on extensive historical data to achieve high prediction accuracy. This dependency can limit their effectiveness in scenarios with limited data availability, such as newly established systems or rapidly changing environments. These issues motivate the development of novel architectures and improvements in linear models, which aim to address these shortcomings.
GLinear achieves superior results by integrating two components: (1) a non-linear Gaussian Error Linear Unit (GELU)-based transformation layer to capture intricate patterns and (2) Reversible Instance Normalization (RevIN) to standardize data distributions across instances, ensuring consistent performance and adaptability across diverse datasets. This approach provides a more comprehensive and efficient solution for TSF.
@@ III Architecture of Different Linear Models
The input of a time series predictor is the $L$ time steps of past input samples (also referred to as the lookup window or input sequence length). A predictor uses this input to predict $T$ time steps of future values (also referred to as output horizon or prediction length), and $C$ refers to channels. The details of different linear predictors are given below:
@@ III-A Linear Predictor (LTSF-Linear)
The first model is a LTSF-Linear predictor [12] where LTSF stands for long-term TSF. The architecture of the linear predictor can be visualized from Figure 1. This predictor is composed of a single fully connected linear layer, or dense layer. LTSF-Linear predictor does not model any spatial correlations. A single temporal linear layer directly regresses historical time series for future prediction via a weighted sum operation, as follows:  [^1] Notation: Scalars are denoted by italic letters, vectors by bold lowercase (e.g., $\mathbf{x}$), and matrices by bold uppercase (e.g., $\mathbf{W}$). The input sequence length is $L$, the prediction length is $T$, the number of channels is $C$, and the batch size is $B$.
$\mathbf{T}=\mathbf{W}\,\mathbf{L}+\mathbf{b},\quad\mathbf{W}\in\mathbb{R}^{T\times L},\;\mathbf{L}\in\mathbb{R}^{L},\;\mathbf{b},\mathbf{T}\in\mathbb{R}^{T}.$ (1)
For simplicity, Eq. (1) is written in the univariate form. In the multivariate case (with $C$ channels) the same linear mapping is applied independently per channel, i.e. for each $c\in\{1,\dots,C\}$ we have $\hat{\mathbf{x}}^{(c)}=\mathbf{W}\,\mathbf{x}^{(c)}+\mathbf{b}$, yielding $\hat{\mathbf{X}}\in\mathbb{R}^{T\times C}$. For batched inputs, $\mathbf{X}\in\mathbb{R}^{B\times L\times C}$ the mapping is applied over the $L$ $\to$ $T$ axis for every $(b,c)$ pair.
@@ III-C DLinear Predictor (Decomposition-based Linear Model)
DLinear [12] is a combination of a decomposition scheme used in Autoformer and FEDformer with linear layers. It first decomposes raw data input into a trend component by a moving average kernel and a remainder (seasonal) component, as shown in Figure 2 (b). Then, two one-layer linear layers are applied to each component, and the two features are summed up to get the final prediction. By explicitly handling trend, DLinear enhances the performance of a vanilla linear model when there is a clear trend in the data.
@@ IV Methodology
Existing linear models use different simple operations like normalization and decomposition. It is worth noting that with the involvement of simpler mathematical operations, it is possible to build powerful linear predictors with some variations of activation functions to extract meaningful results for the required task [27]. Keeping these in mind, a new Gaussian-activated linear GLinear predictor is proposed.
Figure 4: The architecture of the proposed GLinear predictor.
@@
Figure 4 shows the architecture of the GLinear predictor that is composed of two fully connected layers of the same input size, having a GELU nonlinearity in between them. Different configurations of input and layer sizes are tested to lead to this final architecture.
@@
Furthermore, RevIN is applied to the input and output layers of the GLinear model [32]. This normalization layer transforms the original data distribution into a mean-centered distribution, where the distribution discrepancy between different instances is reduced. This normalized data is then applied as new input to the used predictor, and then the final output is de-normalized at the last step to provide the final prediction. RevIN can be used with any existing model. It does not add any significant overhead in training time due to the low computational complexity of this normalization scheme.
@@
The time-series data usually suffers from a distribution shift problem due to changes in its mean and variance over time that can degrade the performance of time-series predictors. RevIN utilizes a learnable affine transformation to remove and restore the statistical information of a time-series instance that can be helpful to handle datasets with distribution shift problems.
IV-A Mathematical Formulation of GLinear
Let the input batch be $\mathbf{X}\in\mathbb{R}^{B\times L\times C}$, where $B$ is the batch size, $L$ is the lookup window (input sequence length), and $C$ is the number of variables (channels). The forecasting objective is to predict $\widehat{\mathbf{Y}}\in\mathbb{R}^{B\times T\times C}$, where $T$ is the output horizon (prediction length).
IV-A1 Reversible Instance Normalization (RevIN)
For each instance $b$ and channel $c$, we compute the per-instance statistics
$\mu_{b,c}\;=\;\frac{1}{L}\sum_{t=1}^{L}X_{b,t,c},\qquad\sigma_{b,c}\;=\;\sqrt{\frac{1}{L}\sum_{t=1}^{L}\big{(}X_{b,t,c}-\mu_{b,c}\big{)}^{2}+\varepsilon}.$ (2)
The normalized (and affine-transformed) input becomes
$\tilde{X}_{b,t,c}\;=\;\gamma_{c}\odot\frac{X_{b,t,c}-\mu_{b,c}}{\sigma_{b,c}}+\beta_{c},\qquad\widetilde{\mathbf{X}}\in\mathbb{R}^{B\times L\times C}.$ (3)
where $\gamma_{c},\beta_{c}$ are learnable per-channel affine parameters. For processing, we reshape each instance/channel to a column vector:
$\mathbf{x}^{(b,c)}=\big{(}\widetilde{X}_{b,1,c},\dots,\widetilde{X}_{b,L,c}\big{)}^{\!\top}\in\mathbb{R}^{L}.$
IV-A2 Gaussian Distribution-based Non-linear Mapping
After reshaping, each channel $c$ is processed independently (channel-wise mapping). For every batch index $b$ and channel $c$, we apply a compact two-layer MLP with the GELU nonlinearity. Let the (per-instance, per-channel) input column vector be $\mathbf{x}^{(b,c)}\in\mathbb{R}^{L}$, we also fixed $\mathbf{H}=\mathbf{L}$ in the experiments.
$\mathbf{h}^{(b,c)}$ $=\mathrm{GELU}\!\big{(}\mathbf{W}^{(1,c)}\mathbf{x}^{(b,c)}+\mathbf{b}^{(1,c)}\big{)},$ (4)
$\quad\mathbf{W}^{(1,c)}\in\mathbb{R}^{L\times L},\;\mathbf{b}^{(1,c)}\in\mathbb{R}^{L},$
$\mathbf{o}^{(b,c)}$ $=\mathbf{W}^{(2,c)}\mathbf{h}^{(b,c)}+\mathbf{b}^{(2,c)},$ (5)
$\quad\mathbf{W}^{(2,c)}\in\mathbb{R}^{T\times L},\;\mathbf{b}^{(2,c)}\in\mathbb{R}^{T}.$
Here $\mathrm{GELU}$  is the Gaussian Error Linear Unit [33], a smooth non-linear activation function, defined as,
$\mathrm{GELU}(x)=x\cdot\Phi(x),\qquad\texttt{where}\quad\Phi(x)=\tfrac{1}{2}\Big{(}1+\mathrm{erf}\!\big{(}\tfrac{x}{\sqrt{2}}\big{)}\Big{)}.$ (6)
For fast approximation, we adopted GELU using
$\mathrm{GELU}(x)\approx 0.5x\Big{(}1+\tanh\big{(}\sqrt{\tfrac{2}{\pi}}(x+0.044715x^{3})\big{)}\Big{)}.$
IV-A3 Denormalization
After the model outputs $\mathbf{o}^{(b,c)}\in\mathbb{R}^{T}$, we reverse the affine and instance scaling:
$\widehat{Y}_{b,t,c}\;=\;\frac{o^{(b,c)}_{t}-\beta_{c}}{\gamma_{c}}\cdot\sigma_{b,c}+\mu_{b,c},\qquad\widehat{\mathbf{Y}}\in\mathbb{R}^{B\times T\times C}.$ (7)
IV-B Features of GLinear
Some features of the GLinear model are